HW 2-1. For the RLC circuit in HW 1-2, with the voltage source x(t) as the input and the loop current y(t) as the output
Answer #1

Solution:

o applying KvL in the circuils malt) – L digesty – Reyet) - Syltido pulting L=1, R=3, 120.5 alt) - dys – Bylt)– 15444 =0 taki5+ 100 105 10 nits alts XVha Ś Y() -12 - 5- sissez by inverve laplace ylte.de 143-2 ja etezt © alt) = 3/6 (104) X(s) 10 Yle